# Revisiting Convergence of AdaGrad with Relaxed Assumptions

### v1

@@ Appendix C Proof of convergence under the generalized smoothness
In this section, we provide the detailed proof of Theorem 2. We still follow all the notations defined in Section 5 and $\mathcal{G}_{s}$ in (9). We shall first introduce a sequence $\{\mathcal{L}_{s}\}_{s\geq 1}$ as follows:
$\mathcal{L}_{s}=2L_{0}+2L_{1}\left(4L_{1}\Delta^{(x)}_{s}+\sqrt{4L_{0}\Delta^{(x)}_{s}}\right).$ (44)
@@ C.1 Convergence of AdaGrad
Let $T\geq 1$ and $\delta\in(0,1)$. Suppose that $\{\bm{x}_{s}\}_{s\in[T]}$ is a sequence generated by Algorithm 1 with $\beta=0$, $f$ is $(L_{0},L_{1})$-smooth satisfying (25), Assumptions (A1), (A2) hold and Assumption (A3) holds with $A=0$, and the parameters follow the condition in (27) with $\beta=0$, $\mathcal{H},\mathcal{L}$ follow the definitions in (28),
@@ C.2 Technical lemmas
[…]
@@ C.3 Rough Estimations
Since (47) holds, we could rely on (52) and obtain that
@@
Finally, we sum up over $t\in[T]$ and obtain the desired result. ∎
@@ C.4 Start Point and Decomposition
where we use $\bm{x}_{1}=\bm{y}_{1}$ and the definition of A in (10). We also follow the same decomposition for A in (11) and (13).
@@ C.5 Estimating A
The technique for estimating A.1.1 is similar to Lemma 5.3. However, we shall take a different setting of $\lambda=1/(3\mathcal{H})$ in (34) where $\mathcal{H}$ is as in (28), leading to the new estimation as follows: with probability at least $1-\delta$,
${\bf A.1.1}\leq\frac{1}{4}\sum_{s=1}^{t}\frac{\mathcal{G}_{s}}{\mathcal{H}}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}+3\mathcal{H}\log\left(\frac{T}{\delta}\right),\quad\forall t\in[T].$ (58)
The estimation for A.1.2 in (15) remains unchanged since it does not rely on smoothness. Finally, the estimation for A.2 should be replaced by the following lemma.
@@ C.6 Bounding the function value gap
$\Delta^{(x)}_{t}\leq\Lambda_{x},\quad\mathcal{G}_{t}\leq\mathcal{H},\quad\mathcal{L}_{t}\leq\mathcal{L},\quad\forall t\in[T+1],$ (61)
@@
$\Delta^{(y)}_{t+1}\leq\Lambda_{y}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2},\quad\forall t\in[T],$ (62)
where $\Lambda_{x},\Lambda_{y},\mathcal{H},\mathcal{L}$ are as in Theorem 2.
@@
Plugging the estimations (58) for A.1.1, (15) for A.1.2, (59) for A.2, and (13) into (11), and then combining with (57), and subtracting $f^{*}$ on both sides,
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}+\frac{\eta}{1-\beta}\sum_{s=1}^{t}\left(\frac{\mathcal{G}_{s}}{4\mathcal{H}}-\frac{3}{4}\right)\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}+\frac{3\mathcal{H}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)+\frac{\eta}{1-\beta}\sum_{s=1}^{t}\mathcal{G}_{s}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}$
@@
We still rely on an induction argument to deduce the result. First, we define two polynomials ~$\tilde{\mathcal{I}}_{t}$ and $\tilde{\mathcal{J}}_{t}$ with respect to $t$ and two values $\Lambda_{y}$ and $\Lambda_{x}$ that are independent from $t$ as follows:
@@
It’s worthy noting that ~$\tilde{\mathcal{I}}_{t}$ and $\tilde{\mathcal{J}}_{t}$ are deterministic polynomials with respect to $t$ and are dependent on hyper-parameters $C_{0},\beta,d$ and problem-parameters $A,B,C$. We could verify that $\Lambda_{y}\sim\mathcal{O}(\text{poly}(\log(T/\delta)))$ and $\Lambda_{x}\sim\mathcal{O}(\text{poly}(\log(T/\delta)))$.
@@
$\Delta^{(x)}_{s}\leq\Lambda_{x},\quad\text{thus},\quad\mathcal{G}_{s}\leq\mathcal{H},\quad\mathcal{L}_{s}\leq\mathcal{L},\quad\forall s\in[t],$ (68)
where we rely on $\mathcal{G}_{s}$ in (9), $\mathcal{L}_{s}$ in (44), and $\mathcal{H},\mathcal{L}$ in (28). We thus apply (68) to (C.6), and we get
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}+\frac{3\mathcal{H}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)+\frac{\eta\mathcal{H}}{1-\beta}\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}$
@@
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}+\frac{3\mathcal{H}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)+\frac{\mathcal{H}\eta d}{1-\beta}\log\mathcal{J}_{t}$
@@
Combining with (27), we have that
@@
$\Delta^{(y)}_{t+1}\leq\Lambda_{y}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}\leq\Lambda_{y}.$ (74)
@@
$\Delta^{(x)}_{s}$ $\leq\Delta^{(y)}_{s}+\frac{1}{2}\|\nabla f(\bm{y}_{s})\|^{2}+\frac{(\mathcal{L}_{s}+1)\beta^{2}}{2(1-\beta)^{2}}\|\bm{x}_{s}-\bm{x}_{s-1}\|^{2}$
$\leq\Delta^{(y)}_{s}+8L_{1}\left(\Delta^{(y)}_{s}\right)^{2}+2L_{0}\Delta^{(y)}_{s}+\frac{\mathcal{L}_{s}+1}{2(1-\beta)^{2}}\|\bm{m}_{s-1}\|^{2},$
@@
where the last inequality applies (72) and (73). Recalling $\Lambda_{x}$ in (67), we found that
@@ C.7 Proof of the main result
$\frac{\eta}{2(1-\beta)}\sum_{s=1}^{T}\frac{\left\|\bar{\bm{g}}_{s}\right\|^{2}}{\|\bm{a}_{s}\|_{\infty}}\leq\frac{\eta}{2(1-\beta)}\sum_{s=1}^{T}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\bm{a}_{s}}}\right\|^{2}\leq\Lambda_{y}.$ (75)
In what follows, we assume that (61) and (62) always hold and deduce the convergence bound under these two inequalities. Note that (61) and (62) hold with probability at least $1-\delta$ according to Proposition C.1. Thus, the final convergence bound also holds with probability at least $1-\delta$. Applying $\bm{a}_{s}$ in (12) and (61), and following the similar analysis in (5.6),
$\|\bm{a}_{s}\|_{\infty}-\epsilon$ $\leq\sqrt{2\sum_{j=1}^{s-1}(A\Delta^{(x)}_{j}+(B+1)\|\bar{\bm{g}}_{j}\|^{2}+C)+\mathcal{G}_{s}^{2}}\leq\sqrt{2(B+1)\sum_{j=1}^{s-1}\|\bar{\bm{g}}_{j}\|^{2}+2(A\Lambda_{x}+C)s+\mathcal{H}^{2}}$

### v2

@@ Appendix C Proof of convergence under the generalized smoothness
In this section, we provide the detailed proof of Theorem 2. We still follow all the notations defined in Section 5. We shall first introduce two sequences $\{\mathcal{H}_{s}\}_{s\geq 1}$ and $\{\mathcal{L}_{s}\}_{s\geq 1}$ as follows:
$\mathcal{H}_{s}=\sqrt{2A\Delta^{(x)}_{s}+2(B+1)\left(4L_{1}\Delta^{(x)}_{s}+\sqrt{4L_{0}\Delta^{(x)}_{s}}\right)^{2}+2C},\quad\mathcal{L}_{s}=2L_{0}+2L_{1}\left(4L_{1}\Delta^{(x)}_{s}+\sqrt{4L_{0}\Delta^{(x)}_{s}}\right).$ (45)
As a consequence, we slightly change the proxy step-size $\bm{a}_{s}$ in (12) as
$\tilde{\bm{a}}_{s}=\sqrt{\bm{v}_{s-1}+\left(\mathcal{H}_{s}{\bf 1}_{d}\right)^{2}}+\bm{\epsilon},\quad\forall s\in[T].$ (46)
@@ C.1 Convergence of AdaGrad
Let $T\geq 1$ and $\delta\in(0,1)$. Suppose that $\{\bm{x}_{s}\}_{s\in[T]}$ is a sequence generated by Algorithm 1 with $\beta=0$, $f$ is $(L_{0},L_{1})$-smooth satisfying (25), Assumptions (A1), (A2) hold and Assumption (A3) holds with $A=0$, and the hyper-parameters follow the condition in (27) with $\beta=0$, $\mathcal{H},\mathcal{L}$ follow the definitions in (28),
@@ C.2 Technical lemmas
$\eta\leq\frac{(1-\beta)^{2}}{L_{1}\sqrt{d}},\quad\text{then},\quad\max\{\|\bm{x}_{s+1}-\bm{x}_{s}\|,\|\bm{y}_{s}-\bm{x}_{s}\|,\|\bm{y}_{s+1}-\bm{y}_{s}\|\}\leq\frac{1}{L_{1}},\quad\forall s\geq 1.$ (49)
@@ C.3 Rough Estimations
Since (49) holds, we could rely on the updated rule in Algorithm 1 and (54) to obtain that
@@
We define $\sum_{a}^{b}=0$ when $a<b$. Then, we sum up over $t\in[0,1,\cdots T-1]$ and obtain the desired result. ∎
@@ C.4 Start Point and Decomposition
where we use $\bm{x}_{1}=\bm{y}_{1}$ and the definition of A in (10). We follow the decomposition in (60) and restate as follows,
@@ C.5 Estimating A
We then introduce $\tilde{\bm{a}}$ in (46) into (60) to derive that
${\bf A.1}=-\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2}\underbrace{-\sum_{s=1}^{t}\left\langle\bar{\bm{g}}_{s},\frac{\bm{\xi}_{s}}{\tilde{\bm{a}}_{s}}\right\rangle}_{\textbf{A'.1.1}}+\underbrace{\sum_{s=1}^{t}\left\langle\bar{\bm{g}}_{s},\left(\frac{1}{\tilde{\bm{a}}_{s}}-\frac{1}{\bm{b}_{s}}\right)\bm{g}_{s}\right\rangle}_{\textbf{A'.1.2}}.$ (61)
The technique for estimating A’.1.1 is similar to Lemma 5.3. Let $X^{\prime}_{s}=-\left\langle\bar{\bm{g}}_{s},\frac{\bm{\xi}_{s}}{\tilde{\bm{a}}_{s}}\right\rangle$. We could still verify that $X^{\prime}_{s}$ is a martingale difference sequence and define $\zeta^{\prime}_{s}=\left\|{\bar{\bm{g}}_{s}\over\tilde{\bm{a}}_{s}}\right\|\sqrt{A\Delta^{(x)}_{s}+B\|\bar{\bm{g}}_{s}\|^{2}+C}$. Similarly, $\zeta^{\prime}_{s}$ is a random variable dependent on $\bm{z}_{1},\cdots,\bm{z}_{s-1}$. Using Cauchy-Schwarz inequality and Assumption (A3), we have
$\mathbb{E}\left[\exp\left[\left(\frac{X^{\prime}_{s}}{\zeta^{\prime}_{s}}\right)^{2}\right]\mid\bm{z}_{1},\cdots,\bm{z}_{s-1}\right]\leq\mathbb{E}\left[\exp\left(\frac{\|\bm{\xi}_{s}\|^{2}}{A\Delta^{(x)}_{s}+B\|\bar{\bm{g}}_{s}\|^{2}+C}\right)\mid\bm{z}_{1},\cdots,\bm{z}_{s-1}\right]\leq\mathrm{e}.$
However, using Lemma C.1 and $\mathcal{H}_{s}$ in (45), we derive that
$A\Delta^{(x)}_{s}+B\|\bar{\bm{g}}_{s}\|^{2}+C\leq A\Delta^{(x)}_{s}+B\left(4L_{1}\Delta^{(x)}_{s}+\sqrt{4L_{0}\Delta^{(x)}_{s}}\right)^{2}+C\leq\mathcal{H}_{s}^{2}.$
Hence, we obtain a different inequality from (34) that
A’.1.1 $\leq\frac{3\lambda}{4}\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\bar{g}_{s,i}^{2}}{\tilde{a}_{s,i}^{2}}\left(A\Delta^{(x)}_{s}+B\|\bar{\bm{g}}_{s}\|^{2}+C\right)+\frac{1}{\lambda}\log\left(\frac{1}{\delta}\right)$
$\leq\frac{3\lambda}{4}\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\bar{g}_{s,i}^{2}\mathcal{H}_{s}^{2}}{\tilde{a}_{s,i}^{2}}+\frac{1}{\lambda}\log\left(\frac{1}{\delta}\right)\leq\frac{3\lambda}{4}\sum_{s=1}^{t}\sum_{i=1}^{d}\frac{\bar{g}_{s,i}^{2}\mathcal{H}_{s}}{\tilde{a}_{s,i}}+\frac{1}{\lambda}\log\left(\frac{1}{\delta}\right),$ (62)
where the last inequality applies $1/\tilde{a}_{s,i}\leq 1/\mathcal{H}_{s}$ from (46). Then, we can re-scale $\delta$ and take $\lambda=1/(3\mathcal{H})$, leading to the new estimation as follows: with probability at least $1-\delta$,
[…]
The estimation for A’.1.2 remains similar to (15). We first derive from the basic inequality and Assumption (A3) that
$\|\bm{g}_{s}\|^{2}\leq 2\|\bar{\bm{g}}_{s}\|^{2}+2\|\bm{\xi}_{s}\|^{2}\leq 2A\Delta^{(x)}_{s}+2(B+1)\|\bar{\bm{g}}_{s}\|^{2}+2C\leq\mathcal{H}_{s}^{2}.$
Then, based on $\|\bm{g}_{s}\|^{2}\leq\mathcal{H}_{s}^{2}$, we derive a similar result as in (36) where
$\left|\frac{1}{\tilde{a}_{s,i}}-\frac{1}{b_{s,i}}\right|\leq\frac{\mathcal{H}_{s}}{\tilde{a}_{s,i}b_{s,i}},\quad\forall s\in[T],\forall i\in[d].$
Then, using a similar deduction in (15), we derive that
A’.1.2 $\leq\frac{1}{4}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2}+\sum_{s=1}^{t}\mathcal{H}_{s}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}.$ (64)
The estimation for A.2 in (60) is revised by the following lemma.
@@ C.6 Bounding the function value gap
$\Delta^{(x)}_{t}\leq\Lambda_{x},\quad\mathcal{H}_{t}\leq\mathcal{H},\quad\mathcal{L}_{t}\leq\mathcal{L},\quad\forall t\in[T+1],$ (67)
@@
$\Delta^{(y)}_{t+1}\leq\Lambda_{y}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2},\quad\forall t\in[T],$ (68)
where $\mathcal{H}_{t},\mathcal{L}_{t}$ are as in (45) and $\Lambda_{x},\Lambda_{y},\mathcal{H},\mathcal{L}$ are as in Theorem 2.
@@
Plugging the estimations (63) and (64) into (61), we obtain the bound for A.1. Then, combining with (65), (60) and (59), and subtracting $f^{*}$ on both sides of (59),
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}+\frac{\eta}{1-\beta}\sum_{s=1}^{t}\left(\frac{\mathcal{H}_{s}}{4\mathcal{H}}-\frac{3}{4}\right)\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2}+\frac{3\mathcal{H}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)+\frac{\eta}{1-\beta}\sum_{s=1}^{t}\mathcal{H}_{s}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}$
@@
We still rely on an induction argument to deduce the result. First, we define two polynomials ~$\tilde{\mathcal{I}}_{t}$ and $\tilde{\mathcal{J}}_{t}$ with respect to $t$ and present the detailed expressions of $\Lambda_{y}$ and $\Lambda_{x}$ that are independent from $t$ as follows:
@@
It’s worthy noting that ~$\tilde{\mathcal{I}}_{t}$ and $\tilde{\mathcal{J}}_{t}$ are deterministic polynomials with respect to $t$ and are dependent on hyper-parameters $C_{0},\beta,d$ and problem-parameters $A,B,C$. We could verify that $\Lambda_{y}\sim\mathcal{O}(\log(T/\delta)))$ and $\Lambda_{x}\sim\mathcal{O}(\log^{2}(T/\delta))$.
@@
$\Delta^{(x)}_{s}\leq\Lambda_{x},\quad\text{thus},\quad\mathcal{H}_{s}\leq\mathcal{H},\quad\mathcal{L}_{s}\leq\mathcal{L},\quad\forall s\in[t],$ (74)
where we rely on $\mathcal{H}_{s}$, $\mathcal{L}_{s}$ in (45), and $\mathcal{H},\mathcal{L}$ in (28). We thus apply (74) to (C.6) and get
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2}+\frac{3\mathcal{H}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)+\frac{\eta\mathcal{H}}{1-\beta}\sum_{s=1}^{t}\left\|\frac{\bm{g}_{s}}{\bm{b}_{s}}\right\|^{2}$
@@
$\Delta^{(y)}_{t+1}$ $\leq\Delta^{(x)}_{1}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2}+\frac{3\mathcal{H}\eta}{1-\beta}\log\left(\frac{T}{\delta}\right)+\frac{\mathcal{H}\eta d}{1-\beta}\log\mathcal{J}_{t}$
@@
Using (27), we have that
@@
$\Delta^{(y)}_{t+1}\leq\Lambda_{y}-\frac{\eta}{2(1-\beta)}\sum_{s=1}^{t}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2}\leq\Lambda_{y}.$ (80)
@@
$\Delta^{(x)}_{s}$ $\leq\Delta^{(y)}_{s}+\frac{1}{2}\|\nabla f(\bm{y}_{s})\|^{2}+\frac{(\mathcal{L}_{s}+1)\beta^{2}}{2(1-\beta)^{2}}\|\bm{x}_{s}-\bm{x}_{s-1}\|^{2}\leq\Delta^{(y)}_{s}+8L_{1}\left(\Delta^{(y)}_{s}\right)^{2}+2L_{0}\Delta^{(y)}_{s}+\frac{\mathcal{L}_{s}+1}{2(1-\beta)^{2}}\|\bm{m}_{s-1}\|^{2},$
@@
where the last inequality applies (78) and (79). Recalling $\Lambda_{x}$ in (73), we find that
@@ C.7 Proof of the main result
$\frac{\eta}{2(1-\beta)}\sum_{s=1}^{T}\frac{\left\|\bar{\bm{g}}_{s}\right\|^{2}}{\|\tilde{\bm{a}}_{s}\|_{\infty}}\leq\frac{\eta}{2(1-\beta)}\sum_{s=1}^{T}\left\|\frac{\bar{\bm{g}}_{s}}{\sqrt{\tilde{\bm{a}}_{s}}}\right\|^{2}\leq\Lambda_{y}.$ (81)
In what follows, we assume that (67) and (68) always hold and deduce the convergence bound under these two inequalities. Note that (67) and (68) hold with probability at least $1-\delta$ according to Proposition C.1. Thus, the final convergence bound also holds with probability at least $1-\delta$. Applying $\tilde{\bm{a}}_{s}$ in (46) and (67), and following the similar analysis in (5.6),
$\|\tilde{\bm{a}}_{s}\|_{\infty}-\epsilon$ $\leq\sqrt{2\sum_{j=1}^{s-1}(A\Delta^{(x)}_{j}+(B+1)\|\bar{\bm{g}}_{j}\|^{2}+C)+\mathcal{H}_{s}^{2}}\leq\sqrt{2(B+1)\sum_{j=1}^{s-1}\|\bar{\bm{g}}_{j}\|^{2}+2(A\Lambda_{x}+C)s+\mathcal{H}^{2}}$
@@
Appendix D Experiment
In this section, we present an experiment to verify that AdaGrad can find a stationary point when the noise satisfies Assumption (A3).
Figure 1: Training loss vs. steps and gradient norms vs. steps using SGD and AdaGrad.
Experimental Setup
We follow the experimental task outlined in [Khaled and Richtárik, 2023], where the objective is to minimize a regularized logistic regression problem defined as follows:
$\min_{x\in\mathbb{R}^{d}}\left\{\frac{1}{n}\sum_{i=1}^{n}\log\left(1+\exp(-a_{i}^{\top}x)\right)+\lambda\sum_{j=1}^{d}\frac{x_{j}^{2}}{1+x_{j}^{2}}\right\}.$ (82)
In [Khaled and Richtárik, 2023], it was verified that using uniform sampling over the a9a dataset and the loss function in (82), the noise conforms to Assumption (A3) with $\hat{A}=10.09,\hat{B}=0,\hat{C}=0.373$, which closely aligns with the theoretical values where $A=9,B=0,C=0.994$. We then executed both SGD and AdaGrad to minimize (82) using a batch size of 256. The learning rates were set to $7\times 10^{-6}$ for SGD and $5\times 10^{-4}$ for AdaGrad.
We utilized the a9a dataset and the PyTorch implementations of SGD and AdaGrad. The experiments were conducted on a single NVIDIA GeForce RTX 4090 GPU.
Results
We plotted the training loss and gradient norms against the number of steps in Figure 1, training for 1200 epochs. The results indicate that both SGD and AdaGrad can find a stationary point given a sufficiently large number of steps $T$, thereby supporting the conclusion in Theorem 1.